The disposable income of Mehta family increases from Rs 5000 to Rs 15,000. As a result, the family‘s demand for milk and milk goods has increased from 30 liters to 60 liters per month. Calculate the income elasticity of demand.



1
Expert's answer
2021-07-08T15:19:36-0400
"E=\frac{Q_2-Q_1}{I_2-I_1}\times\frac{I_2+I_1}{Q_2+Q_1}=0.67"